OpenStudy (anonymous):

In 1996, the amount of waste discarded was 15.3 million tons, by 1998 the amount declined to 13.2 million tons. Assume the amount of waste is decreasing according to the exponental decay model

OpenStudy (amistre64):

exponential eh.....

OpenStudy (amistre64):

so we have a time span of 2; a beginning population of 15.3 and an ending population of 13.2

OpenStudy (anonymous):

i got all the way to the end w(9)=15.2e^-0.0706(9) but i dont know what to do next

OpenStudy (amistre64):

13.2 = 15.3 e^2k 13.2/15.3 = e^2k ln(13.2/15.3) = 2k ln(13.2/15.3) ----------- = k =abt. -.0738 2

OpenStudy (anonymous):

sry thats 15.2

OpenStudy (amistre64):

.... hmm. same setup and following; but no time to change it. Me doors are closing

OpenStudy (amistre64):

the k is the constant of variation and is used inthe equation to make it flow right

OpenStudy (amistre64):

y = starting population * e^k

OpenStudy (amistre64):

or perhaps thats e^2k ? prolly

OpenStudy (anonymous):

e^-2k

OpenStudy (amistre64):

gotta run; good luck ;)

OpenStudy (amistre64):

if you get back to this; the 'k' that I was looking for is the constant part of the equation and once determined, the rest can be stripped away .... well, almost all the rest :) still need a principal to begin with

OpenStudy (amistre64):

Now = Then * e^(timespan*krate) 13.2 = 15.2 e^2k ; this is the info they give us ... if I read you right; now /15.2 13.2/15.2 = e^2k ; now we log the sides to drop down our exponent ln(13.2/15.2) = 2k ; now we /2 to get k all alone ln(13.2/15.2) ----------- = k ; and this is the gold at the end of the rainbow. 2 k = -0.0705.... round it to whatever they want to round it to. Now repaint the equation with this value. Future = Start * e^(-.0705*timespan) F = 15.2 * e^(-.0705t)